OpenStudy (anonymous):

Find the height of the triangle. A. 3.4 B. 9.4 C. 3.6 D. 6.6

OpenStudy (anonymous):



OpenStudy (anonymous):

"X"=3.4 as sin"Y"=(opposite)/(Hypotenys) then "X" =sin"20" times 10

OpenStudy (anonymous):

So.. What do i do?

OpenStudy (anonymous):

you need the base and the height also a calculator

OpenStudy (anonymous):

It asks to find the height..

OpenStudy (anonymous):

oh i thought you needed the area, you just need the height \(x\) now \(\sin(20)=\frac{x}{10}\) so \[x=10\sin(20)\] and a calculator gives you the answer

OpenStudy (anonymous):

make sure you calculator is in "degree" mode let me know what you get, and i will check it

OpenStudy (anonymous):

Okay. i got it. thank you.